I really need help. I have only 1 question. here is is the info.
You are driving from Grand Rapids, to Detroit, MI. You leave Grand Rapid at 4:00pm. At 5:10pm you pass through Lansing, MI a distance of 65miles.
I have to write a linear equation that gives the distance in miles, d, in terms of time, t. Let t represents the number of the minutes since 4:00pm. Then, what time will you arrive in Detroit if it is 150 miles from Grand Rapidsm MI.
Okay, I think: d = 65 miles, and t = 70 minutes.
65 divided by 70 = 0.928 miles per minute.
0.928 x 150 minutes = 2.33 hours (arrival time would be at 6 hours and 33 minutes.
Thank you for any help that you can give me.

Answer by stanbon(75887)

You are driving from Grand Rapids, to Detroit, MI.
You leave Grand Rapid at 4:00pm. At 5:10pm you pass through Lansing, MI a distance of 65miles.
---
d = rt
---
t = 70 min = 7/6 hr
d = 65
Rate = 65/(70) miles/min
---
distance = 0.928571t

I have to write a linear equation that gives the distance in miles, d, in terms of time, t.
------
Let t represents the number of the minutes since 4:00pm. Then, what time will you arrive in Detroit if it is 150 miles from Grand Rapidsm MI.
---
150 = 0.028571t
t = 161.54 minutes = (161.54) = (120 + 41.54)min = 2 hr 41.54 min
---
time of arrival at Grand Rapids = 6:54 PM
===============================
Cheers,
Stan H.
========